OpenStudy (anonymous):

the equation r=3cos(theta) represents a circle. Find the coordinates for the center and the length of the radius

OpenStudy (anonymous):

r=\[r=3\cos(\theta)\]

terenzreignz (terenzreignz):

Well, you could try to convert everything to rectangular coordinates first.

OpenStudy (anonymous):

I tried doing it, but I got confused. Here, I'll show you what I did, just a sec

OpenStudy (anonymous):

I converted it to \[x^2+y^2=9x^2\] is that first step right?

terenzreignz (terenzreignz):

How?

OpenStudy (anonymous):

My professor told us that r=x^2 + y^2 and that x=rcos(theta)

OpenStudy (anonymous):

he has that square root as equaling r, and r^2=x^2 +y^2, are you positive?

terenzreignz (terenzreignz):

If you're confused, let's look at polar coordinates again... 

terenzreignz (terenzreignz):